OpenStudy (anonymous):

the table shows that the amount of milk that wisconsin dairy farms produced from 1939 to 1962 which linear best fits data? year milk produced in billions of lbs 1939 12 1962 15 2005 21

OpenStudy (anonymous):

still need help

OpenStudy (anonymous):

1939 would be 15 62 is 12 and 2005 would be 21

OpenStudy (anonymous):

Can someone please explain this problem to me

OpenStudy (anonymous):

you didnt list any linear regressions to pick from? there are no choices

OpenStudy (anonymous):

That is what the question asks. possible answer choices are; y=0.142x-6.38 y=0.142+6.38 y=6.38x+o.142 y=-0.142x+6,38

OpenStudy (anonymous):

try a linear regression with the points (39, 12) (62, 15) (105, 21) do you have statdisk or statcrunch or something that generates linear regression?

OpenStudy (anonymous):

are you allowed to use a calculator that makes a regression??